OpenStudy (anonymous):

Solve the equation: -3 2/5 +b=8 1/5 A. b= 11 3/5 B. b= 4 4/5 C. b= 5 3/5 D. b= 11 1/5 *The Spaces indicate its a Whole number and a fraction for Ex. 15 3/5

OpenStudy (texaschic101):

-3 2/5 + b = 8 1/5 - 17/5 + b = 41/5 add 17/5 to both sides b = 41/5 + 17/5 can you finish ?

OpenStudy (anonymous):

d

OpenStudy (texaschic101):

not quite...try again

OpenStudy (texaschic101):

what is 41/5 + 17/5 ?

OpenStudy (anonymous):

lol what >

OpenStudy (texaschic101):

41/5 + 17/5 = 58/5 turn to a mixed number = 11 3/5